In 2014, the cost of the most frequently bough consumer goods was 2.364 times higher than in 1982 (the CPI in 2014 was 236.4), and in 2015, the cost of this basket of goods was 2.378 times higher than in 1982 (the CPI in 2015 was 237.8). Using the CPI method, what was the inflation between 2014 and 2015?  Answers should be in percent form and rounded to the nearest tenth, i.e. 5.3%.)

>>
Answer #1

Using the CPI method, what was the inflation between 2014 and 2015?

CPI in 2014 = 236.4

CPI in 2015 = 237.8

Subtract the current year's CPI from the base year's CPI.

ie 237.8-236.4= 1.4

inflation between 2014 and 2015= 1.4/236.4*100 = 0.6%  Round to 1%

Hope you understood.

Thank you.